OpenStudy (anonymous):

Determine angle theta: Cos theta = 0.1392 Tan theta = 0.6249

OpenStudy (shane_b):

Take the inverse of each...your calculator should have a cos^-1 and tan^-1 function.

OpenStudy (anonymous):

Do what?

OpenStudy (anonymous):

Ohh, okay.

OpenStudy (anonymous):

Thank you, How do i do these: Sec 15 degree 30' determine the value

OpenStudy (shane_b):

First, convert 15 degrees and 30' to decimal degrees: \[15deg 30' = 15+30*\frac{1}{60}=15.5 degrees\] Second, secant is equal to:\[sec \theta=\frac{1}{cos\theta}\]So \[sec \theta=\frac{1}{cos(15.5)}\]